# LLMs Do Not See Age: Assessing Demographic Bias in Automated Systematic Review Synthesis

### v1

@@ 1 Introduction
  • Presenting a novel age-stratified dataset of systematic review primary studies, grouped by population (children, adults, older adults), designed to support demographic-specific evaluation of summarisation models in biomedical domains.
@@ 2.1 Automatic Systematic Reviews
Data extraction and synthesis represent the most resource-intensive and error-prone phases within the systematic review workflow, often requiring significant manual effort and rigour. Surveys of professionals involved in systematic reviews reinforce this notion; 45% of respondents in one study of 194 professionals identified data extraction as the most time-consuming stage (Scott2021). Consequently, there has been an increased focus on automating these steps, particularly using LLMs (ge_leveraging_2024; schmidt_data_2023-1). While LLMs have shown promise in biomedical information extraction, their accuracy remains inconsistent across tasks and domains. More broadly, concerns have emerged about their tendency to amplify societal biases, especially in high-stakes fields like medicine. Biases in training data, model design, and linguistic priors can lead to unequal treatment of groups based on attributes such as race (yang_unmasking_2024) and gender (bajaj-etal-2024-evaluating; tang_gendercare_2024), resulting in harms like stereotyping and misassociation (gallegos_bias_2024). Yet demographic attributes like age remain underexplored in bias evaluations, particularly in biomedical synthesis.
Automatic summarisation selects and integrates key content from one or more documents to produce a concise, informative output (nenkova_automatic_2011). This task is increasingly important in the biomedical domain, where the exponential growth of publications makes manual synthesis difficult to scale (pawar_survey_2023). In particular, multi-document summarisation offers a promising way to approximate the narrative synthesis found in systematic reviews. As LLMs are increasingly adopted for this purpose, it becomes essential to examine not only their efficiency, but also how reliably and equitably they preserve critical population-specific information.
@@ 2.2 Bias in Automated Synthesis and Summarisation
Despite the promise of LLMs for multi-document biomedical summarisation, their outputs often reflect biases that can distort or omit critical demographic information. This is especially concerning in systematic reviews, where under-representation or misrepresentation of population groups may have direct clinical and policy implications. Prior work has examined bias in summarisation across domains such as news (steen-markert-2024-bias), opinion generation (huang-etal-2023-examining; huang-etal-2024-bias), and radiology reports (Seyyed-Kalantari2021; doi:10.1148/radiol.232286). However, biomedicine, particularly narrative synthesis in systematic reviews, has received little attention, even though omissions or hallucinations of population-specific information can compromise the validity of clinical evidence.
Fairness and bias have been long-standing concerns in NLP (bolukbasi_man_2016; bender_data_2018; blodgett-etal-2020-language; tang_gendercare_2024), though their definitions are often inconsistent across tasks and domains. We draw on the framework posited by crawford_trouble_2017, which distinguishes between two forms of algorithmic bias: allocative bias, which affects access to resources, and representational bias, which misrepresents or excludes certain groups (sun_mitigating_2019; suresh_framework_2021). This study focuses on representational bias, particularly the loss or distortion of age-related information in generated summaries. We adopt the definition of an unbiased summariser proposed by steen_bias_2024, which emphasises faithful and complete representation of relevant input content.
@@ 3.1 Dataset Construction
Systematic reviews were selected based on the presence of demographic terms (e.g., ’Aged’, ’Older Adult’, ’Adult’, ’Child’) in titles or Medical Subject Headings (MeSH) annotations. Searches were conducted in three open-access biomedical databases: PubMed (ncbi2025), Cochrane (cochrane2025), and Web of Science (clarivate2025), using combinations of ’systematic review’ and demographic keywords. This yielded reviews across medical domains, categorised into three population groups: children, adults, and older adults. Inclusion required that reviews (i) focus on a single, well-defined demographic group and (ii) cite at least three primary studies with accessible abstracts. Non-compliant reviews were excluded. For each included review, cited primary studies were retrieved via PubMed and PubMed Central identifiers (PMIDs/PMCIDs); if unavailable, in-text hyperlinks were used. The final dataset qualifies under the ukcdpa1988_29a exemption for non-commercial text and data analysis [^1] We release the PubMed IDs for systematic reviews in DemogSummary alongside the code at: https://github.com/Favour-Yahdii/lllms_dont_see_age to support transparency, reproducibility, and further research..
@@ 3.2 Demographic Annotation
We identified demographic information using a combination of rule-based methods and LLM-based named entity recognition, details are included in Appendix B.
@@
Public Health 10
Mental Health 2
Nutrition and Digestive Health 12
Cognitive Health 3
@@ 4.1 Model Selection
GPT:
(OpenAI/gpt-4.1, Nano variant) (openai2023gpt4): a proprietary model accessed via the OpenAI API. We included this model as a strong commercial baseline, selected for its favourable trade-off between speed and performance in general-purpose language understanding tasks.
Longformer:
(allenai/led-large-16384-arxiv) (beltagy2020longformerlongdocumenttransformer): a transformer-based encoder-decoder model designed specifically for long-document processing. It supports input sequences of up to 16,384 tokens and was pre-trained on scientific papers from ArXiv, making it particularly suitable for tasks involving structured academic language and domain-specific content.
@@ 4.2 Task Setup
Each model is tasked with generating an abstractive narrative synthesis based on the full set of primary study abstracts from a systematic review. The target output is a concise summary approximating the published systematic review abstract. We evaluated two prompt conditions: a regular prompt (Table 3), which made no reference to patient demographics, and an age-aware prompt (Table 4), which explicitly informed the model of the population group involved. This setup enabled assessment of whether demographic cues influence summarisation behaviour.
@@ 4.3 Implementation Details
All experiments were conducted over approximately 90 GPU hours on a single NVIDIA L4. Inference with GPT-4.1 (Nano variant) via the OpenAI API cost approximately \$5 in total. The open-source models (QWEN and Longformer) were run using Hugging Face Inference Endpoints, with an estimated total cost of \$4. More details, including runtime environment, inference cost, and hyper parameters, are provided in Appendix  A.
You are an experienced and objective biomedical systematic reviewer. Your task is to draft a concise, structured abstract that approximates a systematic review abstract, using the set of provided biomedical research abstracts. Ensure to produce your summary abstract based only on the provided abstracts. Do not include any external information or personal opinions. Your summary should be a synthesis of the provided abstracts, not a critique or evaluation of them.
Table 3: Regular prompt used for summary generation.
@@
Table 4: Age-aware prompt used for summary generation.
@@ 5.2 Demographic Salience Score (DSS)
Each primary study was subsequently categorised into one of the following age groups:
  • Childhood: <18 years
  • Adulthood: 18–59 years
  • Older Adult: 60+ years
@@
The DSS has two main components, the Entity Retention Score (ERS) and the Hallucination Penalty (HP). Components such as the omission rate are derivatives of these main components. Let $Ent_{summary}$ be the set of demographic entities in the generated summary. We define:
@@
the proportion of entities in $Ent_{summary}$ that do not match any entity in $Ent_{gold}$. A generated entity is considered a match if it either exactly matches a gold entity or exceeds a specified cosine similarity threshold.
@@
where $T_{\text{gen}}$ is the number of generated tokens, and $T_{\text{max}}$ is a predefined token limit.
The adjusted hallucination penalty is then defined as:
$HP^{\prime}=HP+OP$ (5)
Final Score:
@@
$\text{DSS}=\alpha\times\text{ERS}-\gamma\times\text{HP}^{\prime}$ (4)
Where $\alpha$, and $\gamma$ are non-negative weighting coefficients. To normalise the score to the $[0,1]$ range, we divide by the maximum achievable score $\alpha\cdot N$ (i.e., a perfect score with full retention and no hallucination penalty), and clip negative values:
$\text{DSS}_{\text{normalised}}=\max\left(0,\frac{\alpha\times\text{ERS}-\gamma\times\text{HP}}{\alpha\times N}\right)$ (5)
Here, $N$ is the number of systematic reviews. This normalisation ensures interpretability while bounding the score, rewarding summaries that retain salient demographic content and penalising unsupported or excessive additions.
@@
We identified matched demographic entities using cosine-based semantic similarity, applying a threshold of 0.7 to determine matches [^3] We also explored prompting an LLM for entity matching but found it inconsistent and less reproducible.. The same similarity threshold (0.7) was applied inversely to identify hallucinations, i.e., entities with similarity below the threshold were considered unsupported. Weighting parameters $\alpha=\gamma=2$ were used in Equations 4 and 5, balancing the contribution of entity retention and hallucination penalties. These settings were chosen to prioritise demographic salience while discouraging unsupported content  [^4] We performed sensitivity analysis tests for these parameters; details can be found in Appendix  C..
@@ 6 Results and Discussions
Group Model BERT $\uparrow$ BART $\downarrow$ BLEU $\uparrow$ FactCC $\uparrow$ ERS $\uparrow$ HP $\downarrow$ Omission $\downarrow$ OP $\downarrow$ DSS $\uparrow$
@@
Older Adults GPT 0.83 -2.00 3.75 0.71 0.92 0.14 0.08 0 0.78
@@
Children GPT 0.84 -2.25 3.04 0.83 0.87 0.17 0.13 0 0.69
QWEN 0.81 -2.31 0.72 0.63 0.89 2.06 0.11 1.32 0( -2.49)
Longformer 0.82 -2.31 2.04 0.59 0.91 0.33 0.09 0 0.58
Older Adults GPT 0.84 -2.00 3.70 0.67 0.85 0.04 0.15 0 0.81
@@
Based on the results in Table  5, BLEU, BERTScore, and BARTScore exhibit minimal variation across models and age groups, indicating a lack of sensitivity to population-specific fidelity and limits their utility in demographic evaluation.
While limited, FactCC reveals greater differentiation. GPT consistently scores higher, with age-aware prompts improving performance in the child group. However, scores remain low for older adults across QWEN and Longformer, despite similar gains, indicating ongoing challenges in summarising this group. The divergence between high FactCC scores and poor demographic fidelity underscores the need for metrics that capture representational accuracy.
@@
While Table 5 summarises model-level averages, it obscures instance-level variability that can affect the reliability of summarisation in practice. Figure  3 shows the distribution of Entity Retention Score (Figure  3(a)) and Hallucination Score (Figure  3(b)) for the adult group under the regular prompt. The ERS distribution highlights GPT’s consistently strong demographic coverage, with minimal variation across reviews. In contrast, Longformer displays a wide spread and lower median, suggesting inconsistent retention of key population information. QWEN achieves comparable retention but at the cost of highly variable hallucination scores, including extreme outliers, indicating instability in content faithfulness. These patterns suggest that even when models appear comparable on average, their behaviour can diverge substantially across instances, a risk especially critical in biomedical settings. Additional distributions for other age groups and prompt conditions are provided in Appendix G, along with a token-level qualitative analysis in Section  6.3.
@@ 6.3 Qualitative Evaluation and Token Analysis
Hallucinated entities, while less frequent overall, often involved tangential or demographically unrelated populations not mentioned in the gold document. Examples include the inclusion of prisoners in the population group of a review on smoking cessation interventions for young adults. Hallucination frequency was higher for QWEN and more prominent under the age-aware prompt, reflecting a tendency to overgenerate plausible but unsupported population descriptors. These patterns mirror the quantitative hallucination scores and indicate model-specific susceptibility to fabrication and generation of unsupported content, particularly when prompted to focus on demographic detail.
Omissions included both general and highly specific age-related terms that were present in the gold documents but not captured in the generated outputs. These, similar to retained entities, ranged from descriptive phrases (e.g., mean age 73.66 $\pm$ 14.67 years, residents in aged care settings) to developmental or life stage terms such as infancy, childhood, and young adulthood. Across prompts, omissions were most pronounced for Longformer, consistent with its lower quantitative entity retention scores. The omission of precise subgroups may reflect limitations in entity grounding, particularly for complex or compound descriptors. There seemed to be no obvious pattern in the models decision on which entities to retain and which to exclude [^5] More details on the effects of the prompt design can be found in Appendix D..
@@ 7 Conclusion
This study systematically evaluates demographic fidelity in LLM-based biomedical summarisation, revealing inconsistencies in population-specific representation. By introducing the Demographic Salience Score (DSS) and the DemogSummary dataset, we provide tools to quantify these disparities and establish a foundation for more equitable and transparent evidence summarisation.
@@ Appendix A Hyperparameters
The hyperparameters in Table  6 were selected to encourage highly deterministic, concise outputs (temperature = 0) and to reduce redundancy. The maximum output token limit (750) was chosen both to provide sufficient token space for generating detailed summaries that approximate the length and content of gold-standard review abstracts, as well as a guide for penalising overly lengthy syntheses that provide weak and shallow aggregations of biomedical systematic review abstracts.

### v2

@@ 1 Introduction
  • Presenting a novel age-stratified dataset of systematic review primary studies (DemogSummary), grouped by population (children, adults, older adults), designed to support demographic-specific evaluation of summarisation models in biomedical domains.
@@ 2.1 Automatic Systematic Reviews
Data extraction and synthesis represent the most resource-intensive and error-prone phases within the systematic review workflow, often requiring significant manual effort and rigour. Surveys of professionals involved in systematic reviews reinforce this notion; 45% of respondents in one study of 194 professionals identified data extraction as the most time-consuming stage (Scott2021). Consequently, there has been an increased focus on automating these steps, particularly using LLMs (ge_leveraging_2024; schmidt_data_2023-1). While LLMs have shown promise in biomedical information extraction, their accuracy remains inconsistent across tasks and domains. More broadly, concerns have emerged about their tendency to amplify societal biases, especially in high-stakes fields like medicine. Biases in training data, model design, and linguistic priors can lead to unequal treatment of groups based on attributes such as race (yang_unmasking_2024) and gender (bajaj-etal-2024-evaluating; tang_gendercare_2024), resulting in harms like stereotyping and misassociation (gallegos_bias_2024). Yet demographic attributes like age remain underexplored in bias evaluations, particularly in biomedical synthesis. Automatic summarisation selects and integrates key content from one or more documents to produce a concise, informative output (nenkova_automatic_2011). This task is increasingly important in the biomedical domain, where the exponential growth of publications makes manual synthesis difficult to scale (pawar_survey_2023). In particular, multi-document summarisation offers a promising way to approximate the narrative synthesis found in systematic reviews. As LLMs are increasingly adopted for this purpose, it becomes essential to examine not only their efficiency, but also how reliably and equitably they preserve critical population-specific information.
@@ 2.2 Bias in Automated Synthesis and Summarisation
Despite the promise of LLMs for multi-document biomedical summarisation, their outputs often reflect biases that can distort or omit critical demographic information. This is especially concerning in systematic reviews, where under-representation or misrepresentation of population groups may have direct clinical and policy implications. Prior work has examined bias in summarisation across domains such as news (steen-markert-2024-bias), opinion generation (huang-etal-2023-examining; huang-etal-2024-bias), and radiology reports (Seyyed-Kalantari2021; doi:10.1148/radiol.232286). However, biomedicine has received little attention, particularly in narrative synthesis of systematic reviews, where omissions or hallucinations of population-specific information can compromise the validity of clinical evidence. Fairness and bias have been long-standing concerns in NLP (bolukbasi_man_2016; bender_data_2018; blodgett-etal-2020-language; tang_gendercare_2024), though their definitions are often inconsistent across tasks and domains. We draw on the framework posited by crawford_trouble_2017, which distinguishes between two forms of algorithmic bias: allocative bias, which affects access to resources, and representational bias, which misrepresents or excludes certain groups (sun_mitigating_2019; suresh_framework_2021). This study focuses on representational bias, particularly the loss or distortion of age-related information in generated summaries. We adopt the definition of an unbiased summariser proposed by steen_bias_2024, which emphasises faithful and complete representation of relevant input content.
@@ 3.1 Dataset Construction
Systematic reviews were selected based on the presence of demographic terms (e.g., ‘Aged‘, ‘Older Adult‘, ‘Adult‘, ‘Child‘) in titles or Medical Subject Headings (MeSH) annotations. Searches were conducted in three open-access biomedical databases: PubMed (ncbi2025), Cochrane (cochrane2025), and Web of Science (clarivate2025), using combinations of ‘systematic review‘ and demographic keywords. This yielded reviews across medical domains, categorised into three population groups: children (aged <18 years), adults (aged 28-59 years), and older adults (60+ years). Inclusion required that reviews (i) focus on a single, well-defined demographic group and (ii) cite at least three primary studies with accessible abstracts. Any reviews that did not comply with (i) and (ii) were excluded. For each included review, cited primary studies were retrieved via PubMed and PubMed Central identifiers (PMIDs/PMCIDs); if unavailable, in-text hyperlinks were used. The final dataset qualifies under the ukcdpa1988_29a exemption for non-commercial text and data analysis [^1] We release the PubMed IDs for systematic reviews in DemogSummary alongside the code at: https://github.com/Favour-Yahdii/lllms_dont_see_age to support transparency, reproducibility, and future research..
@@ 3.2 Demographic Annotation
We identified demographic information using a combination of rule-based methods and LLM-based named entity recognition; details are included in Appendix B.
@@
Public Health 14
Frailty 2
Mental Health 8
Nutrition and Digestive Health 9
Cognitive Health 1
@@ 4.1 Model Selection
GPT
(OpenAI/gpt-4.1-nano) (openai2023gpt4): a proprietary model accessed via the OpenAI API. We included this model as a strong commercial baseline, selected for its favourable trade-off between speed and performance in general-purpose language understanding tasks.
Longformer
(allenai/led-large-16384-arxiv) (beltagy2020longformerlongdocumenttransformer): a transformer-based encoder-decoder model pretrained on scientific papers from ArXiv and designed specifically for scientific long-document processing, making it particularly suitable for tasks involving structured academic language and domain-specific content.
@@ 4.2 Task Setup
Each model is tasked with generating an abstractive narrative synthesis based on the full set of primary study abstracts from a systematic review. The target output is a concise summary approximating the published systematic review abstract. We evaluated two prompt conditions: a regular prompt, which did not refer to patient demographics, and an age-aware prompt, which explicitly informed the model of the population group involved. Both prompts are outlined in (Table 3). This setup enabled assessment of whether demographic cues influence summarisation behaviour.
@@ 4.3 Implementation Details
All experiments were conducted over approximately 90 GPU hours on a single NVIDIA L4. Inference with GPT-4.1-nano via the OpenAI API cost approximately \$5 for all reviews, while the open-source models (QWEN and Longformer) were accessed via Hugging Face. More details, including runtime environment, inference cost, and hyperparameters, are provided in Appendix  A.
@@
Table 3: Prompt used for summary generation. The text in bold was included only in the age-aware prompt.
@@ 5.2 Demographic Salience Score (DSS)
Each primary study was subsequently categorised into one of the following age groups; Child: <18 years, Adult: 18–59 years, Older Adult: 60+ years.
@@
The DSS has two main components, the Entity Retention Score (ERS) and the Hallucination Penalty (HP). Components such as the omission rate are derivatives of these main components. We establish $Ent_{summary}$ as the set of demographic entities in the generated summary.
@@
The proportion of entities in $Ent_{summary}$ that do not match any entity in $Ent_{gold}$. A generated entity is considered a match if it either exactly matches a gold entity or exceeds a specified cosine similarity threshold333see Section 5.2 for details.
@@
where $T_{\text{gen}}$ is the number of generated tokens, and $T_{\text{max}}$ is a predefined token limit. The adjusted hallucination penalty is then defined as:
$HP_{adj}=HP+OP$ (5)
DSS:
@@
$\text{DSS}=\alpha\times\text{ERS}-\gamma\times\text{HP}_{\text{adj}}$ (6)
Where $\alpha$, and $\gamma$ are non-negative weighting coefficients. To normalise the score to the $[0,1]$ range, we divide by the maximum achievable score $\alpha\cdot N$ (i.e., a perfect score with full retention and no hallucination penalty), and clip negative values, where $N$ is the number of systematic reviews:
$\text{DSS}_{\text{normalised}}=\max\left(0,\frac{\alpha\times\text{ERS}-\gamma\times\text{HP}_{\text{adj}}}{\alpha\times N}\right)$ (7)
This normalisation ensures interpretability while bounding the score, rewarding summaries that retain salient demographic content and penalising unsupported or excessive additions.
@@
We identified matched demographic entities using cosine-based semantic similarity, applying a threshold of 0.7 to determine matches [^4] We also explored prompting an LLM for entity matching but found it inconsistent and less reproducible.. The same similarity threshold was applied inversely to identify hallucinations,i.e., entities with similarity below the threshold were considered unsupported. Weighting parameters $\alpha=\gamma=2$ were used in Equations 6 and 7, balancing the contribution of entity retention and hallucination penalties. These settings were chosen to prioritise demographic salience while discouraging unsupported content  [^5] We performed sensitivity analysis tests for these parameters; details can be found in Appendix  C..
@@ 6 Results and Discussions
Group Model BERT $\uparrow$ BART $\uparrow$ BLEU $\uparrow$ FactCC $\uparrow$ ERS $\uparrow$ HP $\downarrow$ Omission $\downarrow$ OP $\downarrow$ DSS $\uparrow$
@@
Adult GPT 0.83 -2.31 2.05 0.82 0.81 0.12 0.19 0 0.69
QWEN 0.81 -2.30 1.76 0.51 0.78 0.74 0.22 1.02 0 (-0.98)
Longformer 0.80 -2.37 2.16 0.60 0.45 0.18 0.50 0.00 0.27
Older Adult GPT 0.83 -2.00 3.75 0.71 0.92 0.14 0.08 0 0.78
@@
Older Adult GPT 0.84 -2.00 3.70 0.67 0.85 0.04 0.15 0 0.81
@@
Based on the results in Table  4, BLEU, BERTScore, and BARTScore exhibit minimal variation across models and age groups, indicating a lack of sensitivity to population-specific fidelity and limits their utility in demographic evaluation. While limited, FactCC reveals greater differentiation. GPT consistently scores higher, with age-aware prompts improving performance in the child group. However, scores remain low for older adults across QWEN and Longformer, despite similar gains, indicating ongoing challenges in summarising this group. The divergence between high FactCC scores and poor demographic fidelity underscores the need for metrics that capture representational accuracy.
@@
While Table 4 summarises model-level averages, it obscures instance-level variability that can affect the reliability of summarisation in practice. Figure  3 shows the distribution of Entity Retention Score (Figure  3(a)) and Hallucination Score (Figure  3(b)) for the adult group under the regular prompt. The ERS distribution highlights GPT’s consistently strong demographic coverage, with minimal variation across reviews. In contrast, Longformer displays a wide spread and lower median, suggesting inconsistent retention of key population information. QWEN achieves comparable retention but at the cost of highly variable hallucination scores, including extreme outliers, indicating instability in content faithfulness. These patterns suggest that even when models appear comparable on average, their behaviour can diverge substantially across instances, a risk especially critical in biomedical settings. Additional distributions for other age groups and prompt conditions are provided in Appendix F, along with a token-level qualitative analysis in Section  6.3.
@@ 6.3 Qualitative Evaluation and Token Analysis
Hallucinated entities, while less frequent overall, often involved tangential or demographically unrelated populations not mentioned in the gold document. Examples include the inclusion of prisoners in the population group of a review on smoking cessation interventions for young adults. Hallucination frequency was higher for QWEN and more prominent under the age-aware prompt, reflecting a tendency to overgenerate plausible but unsupported population descriptors. These patterns mirror the quantitative hallucination scores and indicate model-specific susceptibility to fabrication and generation of unsupported content, particularly when prompted to focus on demographic detail. Omissions included both general and highly specific age-related terms that were present in the gold documents but not captured in the generated outputs. These, similar to retained entities, ranged from descriptive phrases (e.g., mean age 73.66 $\pm$ 14.67 years, residents in aged care settings) to developmental or life stage terms such as infancy, childhood, and young adulthood. Across prompts, omissions were most pronounced for Longformer, consistent with its lower quantitative entity retention scores. The omission of precise subgroups may reflect limitations in entity grounding, particularly for complex or compound descriptors. There seemed to be no obvious pattern in the models decision on which entities to retain and which to exclude.
6.4 Effect of Prompt Design
@@ 7 Conclusion
Our study shows that current LLMs exhibit systematic disparities in age-related information retention in biomedical summaries. Using DemogSummary and the Demographic Salience Score (DSS), we quantify these biases, finding adult summaries particularly error-prone and underrepresented populations more likely to be hallucinated. These results highlight the need for demographic-aware evaluation and fair summarisation pipelines, paving the way for more equitable and transparent biomedical NLP systems.
@@ Appendix A Hyperparameters
The hyperparameters in Table  5 were selected to encourage highly deterministic, concise outputs (temperature = 0) and to reduce redundancy. The maximum output token limit (750) was chosen both to provide sufficient token space for generating detailed summaries that approximate the length and content of gold-standard review abstracts, as well as a guide for penalising overly lengthy syntheses that provide weak and shallow aggregations of biomedical systematic review abstracts.